This is actually rational expressions, multiplying and simplifying them. I know the formula A/B * C/D= AC/BD but I dont think I am using it correctly! The problem is :
Multiply and simplify
x^2 - 36/ x^2 * x^2 - 6x/ x^2+2x-48
Any help would be great! The ^2 means squared and * is the multiplication symbol. Thank you! Thank you!

Answer by stanbon(48558)

You should try to simplify before you multiply.
Factor where you can. For example
x^2-36 = (x+6)(x_6)
x^2-6x = x(x_6)
x^2+2x-48 = (x+8)(x-6)
Make these changes and reduce fractions
where you can. Then multiply.
You should get
[(x+6)(x-6)]/x(x+8)
Cheers,
Stan H.